OpenStudy (anonymous):

What is the domain and range of y = 2(x+3)^2 - 1?

OpenStudy (anonymous):

Domain is what values x can take. Range is what values y can take. The function is a parabola that opens up and is shifted down one unit.

OpenStudy (anonymous):

D: \[(-\infty, \infty)\] R: \[[-1, \infty)\] Is this right? @uhhhhhh

OpenStudy (anonymous):

yes